Solve the compound inequality.
5 > - 2x + 4 or 8 < or = to -2x + 3
I came up with (∞, -5/2] ∪ (-1/2,∞) Is that correct?

Answer by John10(297)

Solve the compound inequality.
5 > - 2x + 4 or 8 < or = to -2x + 3
We solve each inequality
1) 5 > -2x + 4
5 - 4 > -2x + 4 - 4
1 > -2x
x > -1/2 (we have to change the sign of inequality because we divide both sides by negative number)
2) 8 <= -2x + 3
8 - 3 <= -2x + 3 - 3
5 <= -2x
x <= -5/2 (we have to change the sign of inequality because we divide both sides by negative number)
Thus the solution is (- inf. , -5/2)U (-1/2, inf.).YOUR ANSWER IS 100% CORRECT!